You and your friend are pushing a 286 kg cart on a theatre stage in order to move it to the northwest corner. You pushing the cart witha 52-N force that points directly to the west; while your friend a 89-N force pointing north. what is the acceleration of the cart assuming there is no loss due to friction?
Answer #1

52 N force is applying in the direction of west i.e. -X direction
and 89 N is applying in the direction of north i.e. Y direction
hence the net force on the block
= (522 + 892)1/2 = 103.078 N
We know that force = mass*acceleration
103.078 = 286*a
a = 0.36 m/s2